OpenStudy (anonymous):

T= 2π√(ℓ/32) solve for ℓ anybody?

OpenStudy (turingtest):

\[T=2 \pi \sqrt{l/32}\]\[T^2=4 \pi^2(l/32)\]\[8T^2/4\pi^2=l\]

OpenStudy (turingtest):

this server isn't the best to my computer at times...

OpenStudy (turingtest):

8T^2/pi^2=l

OpenStudy (turingtest):

is the last line

OpenStudy (anonymous):

:) thanks! that makes a lot more sense than what i was doing. derp.